# Effective Hamiltonian approach to the Quantum Phase transitions in the extended Jaynes-Cummings model

### 11 Dec 2023 v2

@@ V Conclusion
In conclusion, the quantum phases in the extended Jaynes-Cummings model was investigated in this paper. To simulate the open dynamics of spin, a non-Hermitian effective Hamiltonian $H_{\text{dis}}$ was constructed using complex discretization approximation for the environment. The validity of this approach was confirmed by examining the spin dynamics in single-excitation subspace, which was determined exactly by solving Eq. 4. By exactly diagonalizing $H_{\text{dis}}$, the spin dynamics could be further examined by analyzing the ground state, which revealed eigenvalue with the smallest real part. It was found that the spin dynamics in single-excitation subspace is strongly related to two key properties of the ground state, the energy gap, defined as the negative real part of its eigenvalue, and PMP of ground state on the basis state. When the energy gap is zero, the spin dynamics, measured by the survival probability of initial state $|\psi(0)\rangle=|1\rangle|0\rangle^{\otimes k}$, decays rapidly. Conversely, if the energy gap is finite, two distinct spin dynamics can be observed, determined by PMP. If PMP occurs at $|1\rangle$, the survival probability stabilizes eventually at a finite value, which indicates the localization of excitation in the spin. While if not, a stretched-like dynamics was observed. This picture is a result of the competition between the energy gap, which protects excitation against the spontaneous emission into environment, and the intrinsic dissipation in the ground state function, which does not support PMP at $|1\rangle$.
By detecting the occurrence of energy gap and PMP at basis state $|1\rangle$, the critical point $\eta_{I}$ or $\eta_{II}$ can be identified, which separates the localized phase from the dissipative phase or the intermediated phase respectively. Additionally, both $\eta_{i}$ and $\eta_{II}$ display relevance to the value of $s$. For $0<s<\sim 0.6$, $\eta_{II}$ is greater than $\eta_{I}$, which leads to the identification of the intermediated phase. While for $s>\sim 0.6$, $\eta_{II}$ tends to be equal to $\eta_{I}$.
@@
Finally, our approach would allow for an exact analysis of the effects of finite temperature on the phase transition, which is usually done by perturbational method or high-temperature approximation tamascelli19 . Further research on this topic will be presented in the future.
@@ ACKNOWLEDGEMENTS
H.T.C. acknowledges the support of Natural Science Foundation of Shandong Province under Grant No. ZR2021MA036. Y. A. Yan acknowledges the support of National Natural Science Foundation of China (NSFC) under Grant No. 21973036. M.Q. acknowledges the support of NSFC under Grant No. 11805092 and Natural Science Foundation of Shandong Province under Grant No. ZR2018PA012. X.X.Y. acknowledges the support of NSFC under Grant No. 12175033 and National Key R&\&&D Program of China (No. 2021YFE0193500).
@@
Figure A1: (Color online) The comparative plots for the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ with $|\psi(0)\rangle=|e\rangle|0\rangle^{\otimes k}$ for different $s$ and $\eta$. The solid line depict the exact result, whereas the empty circles represent results obtained by CDA. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
Figure A2: (Color online) The comparative plot between the exact results (solid line) and the numerics (empty circle) obtained by $\widetilde{H}_{\text{dis}}$. The other parameters are same to those in Fig. A1.
Figure A3: (Color online) Plots for CDA approach to the long-term feature of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ when $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$ are used respectively. It is evident that the numerics by $H_{\text{dis}}$ would become convergent to the results of $\widetilde{H}_{\text{dis}}$ when $N$ increases.
Figs. A1 illustrate explicitly evolution of the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ for different values of $s$ and $\eta$. The solid line is decided by solving Eq. (3) with step length $\delta t=10^{-4}$, and the empty circles represent results obtained by CDA. A slight difference can be observed for $s=1,\eta=0.5$. Fig. A2 illustrates the results obtain by $\widetilde{H}_{\text{dis}}$, which shows obvious deviation from the exact results at short term evolution when $\eta=0.5$.
To justify $\widetilde{H}_{\text{dis}}$, the long-term feature of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ is evaluated respectively by $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$, as shown in Fig. A3. Although the convergency of calculation is rapid for $\eta=0.05$, the evaluation for $s=0.1,0.5$ converges slowly to a stable result with the increase of $N$. By the observation of convergence, it is reasonable to consider the results by $\widetilde{H}_{\text{dis}}$ as stable. This conjecture can further verified by Fig. A3 (a), in which both $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$ give the same predication of spin dymamics.
However, it is admitted that the statement above is based solely on the observation in Figs. A3. Therefore, $\widetilde{H}_{\text{dis}}$ can only applicable for $s=1$. For $s=0$ or $0.2$, the calculation shows that $\widetilde{H}_{\text{dis}}$ could not provide more accurate results than those by $H_{\text{dis}}$.
@@ Appendix II Supplementary materials about the quantum phase transition in single-excitation subspace
Figure A4: (Color online) The numerical fitting for the stretched-like evolution, shown in Figs. 5.
@@
Fig. A4 illustrates the fitting for the stretched evolution of survival probability, observed in Fig. 5 for $\eta=0.013,0.026,0.1$. The fitting function is chosen as $B\exp\left(-At^{\beta}\right)$, in which $A,B$ and $\beta$ are real and non-negative. Evidently, $\beta$ has a magnitude of order $10^{-1}$, and thus demonstrates clearly the stretched-like behavior.
In Fig. A5, the relevance of $\eta_{I}$ and $\eta_{II}$ to the computational parameters $R$ and $N$ is presented. For $N$, it is observed that the values of both $\eta_{I}$ and $\eta_{II}$ reach a steady state with the rise of $N$. However, for $s=1$, $\eta_{I}$ and $\eta_{II}$ are observed to merge into a single value close to $0.1$. On the other hand, with the increase of $R$, $\eta_{I}$ and $\eta_{II}$ increase very slowly for $s=0$ and $s=0.2$. For $s=1$, $\eta_{II}$ decreases gradually, while $\eta_{I}$ stays constant. This features is a result of the fact that increase of $R$ can improve the accuracy of computation, but is not favorable for the enhancement of efficiency.
@@ Appendix III Approaching spin dynamics in double-excitation subspace by Gauss quadrature in real space
In Fig. A6, $P_{e}(t)$ is evaluated using Gauss quadrature method discretization . The weight function $w(x)=\left(x/\omega_{c}\right)^{s}e^{-x/\omega_{c}}$ is chosen to construct the chain Hamiltonian and the number of lattice sites is set to $N=200$. It is evident that $P_{e}(t)$ has a similar evolution as the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ in single-excitation subspace. However, in contrast to CDA approach, strong fluctuations due to finite $N$ appears when $t>\sim 2$, which limits the application of Gauss quadrature in the simulation of long-term dynamics in open quantum systems.

### 06 Apr 2024 v3

@@ V Conclusion
In conclusion, the quantum phases in the extended Jaynes-Cummings model was investigated in this paper. To simulate the open dynamics of spin, a non-Hermitian effective Hamiltonian $H_{\text{dis}}$ was constructed using complex discretization approximation for the environment. The validity of this approach was confirmed by examining the spin dynamics in the single-excitation subspace, which was determined exactly by solving Eq. 4. By exactly diagonalizing $H_{\text{dis}}$, the spin dynamics could be further examined by analyzing the ground state, which revealed an eigenvalue with the smallest real part. It was found that the spin dynamics in single-excitation subspace was strongly related to two key properties of the ground state, the energy gap, defined as the negative real part of its eigenvalue, and the PMP of ground state on the basis state. When the energy gap is zero, the spin dynamics, measured by the survival probability of initial state $|\psi(0)\rangle=|1\rangle|0\rangle^{\otimes k}$, decays rapidly. Conversely, if the energy gap is finite, two distinct spin dynamics can be observed, determined by PMP. If PMP occurs at $|1\rangle$, the survival probability stabilizes eventually at a finite value, which indicates the localization of excitation in the spin. If this does not occur, a stretched-like dynamics was observed. This picture is a result of the competition between the energy gap, which protects excitation against the spontaneous emission into the environment, and the intrinsic dissipation in the ground state function, which does not support PMP at $|1\rangle$.
By detecting the occurrence of an energy gap and PMP at basis state $|1\rangle$, the critical point $\eta_{I}$ or $\eta_{II}$ can be identified, which separates the localized phase from the dissipative phase or the intermediated phase respectively. Additionally, both $\eta_{i}$ and $\eta_{II}$ display relevance to the value of $s$. For $0<s<\sim 0.6$, $\eta_{II}$ is greater than $\eta_{I}$, which leads to the identification of the intermediated phase. While for $s>\sim 0.6$, $\eta_{II}$ tends to be equal to $\eta_{I}$.
@@ ACKNOWLEDGEMENTS
H.T.C. acknowledges the support of the Natural Science Foundation of Shandong Province under Grant No. ZR2021MA036. Y. A. Y. acknowledges the support of the National Natural Science Foundation of China (NSFC) under Grant No. 21973036. M.Q. acknowledges the support of NSFC under Grant No. 11805092 and the Natural Science Foundation of Shandong Province under Grant No. ZR2018PA012. X.X.Y. acknowledges the support of NSFC under Grant No. 12175033 and the National Key R&\&&D Program of China (Grant No. 2021YFE0193500).
@@
Figure A1: (Color online) The comparative plots for the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ with $|\psi(0)\rangle=|e\rangle|0\rangle^{\otimes k}$ for different $s$ and $\eta$. The solid line depict the exact result, whereas the empty circles represent results obtained by CDA. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen. The evolution time $t$ is in units of $1/\Delta$.
Figure A2: (Color online) The comparative plot between the exact results (solid line) and the numerics (empty circle) obtained by $\widetilde{H}_{\text{dis}}$. The other parameters are same to those in Fig. A1. The evolution time $t$ is in units of $1/\Delta$.
Figure A3: (Color online) Plots for CDA approach to the long-term feature of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ when $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$ are used respectively. It is evident that the numerics by $H_{\text{dis}}$ would become convergent to the results of $\widetilde{H}_{\text{dis}}$ when $N$ increases. The evolution time $t$ is in units of $1/\Delta$.
Figure A1 illustrates explicitly the evolution of the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ for different values of $s$ and $\eta$. The solid line is decided by solving Eq. (3) with step length $\delta t=10^{-4}$, and the empty circles represent results obtained by CDA. A slight difference can be observed for $s=1,\eta=0.5$. Figure A2 illustrates the results obtain by $\widetilde{H}_{\text{dis}}$, which shows obvious deviation from the exact results at short term evolution when $\eta=0.5$.
To justify $\widetilde{H}_{\text{dis}}$, the long-term feature of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ is evaluated respectively by $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$, as shown in Fig. A3. Although the convergency of the calculation is rapid for $\eta=0.05$, the evaluation for $s=0.1,0.5$ converges slowly to a stable result with the increase of $N$. By the observation of convergence, it is reasonable to consider the results by $\widetilde{H}_{\text{dis}}$ as stable. This conjecture can further be verified by Fig. A3 (a), in which both $\widetilde{H}_{\text{dis}}$ and $H_{\text{dis}}$ give the same predication of spin dymamics.
@@ Appendix II Supplementary materials about the quantum phase transition in single-excitation subspace
Figure A4: (Color online) The numerical fitting for the stretched-like evolution, shown in Figs. 5. The evolution time $t$ is in units of $1/\Delta$.
@@
Figure A4 illustrates the fitting for the stretched evolution of survival probability, observed in Fig. 5 for $\eta=0.013,0.026,0.1$. The fitting function is chosen as $B\exp\left(-At^{\beta}\right)$, in which $A,B$ and $\beta$ are real and non-negative. Evidently, $\beta$ has a magnitude of order $10^{-1}$, and thus demonstrates clearly the stretched-like behavior.
In Fig. A5, the relevance of $\eta_{I}$ and $\eta_{II}$ to the computational parameters $R$ and $N$ is presented. For $N$, it is observed that the values of both $\eta_{I}$ and $\eta_{II}$ reach a steady state with the rise of $N$. However, for $s=1$, $\eta_{I}$ and $\eta_{II}$ are observed to merge into a single value close to $0.1$. On the other hand, with the increase of $R$, $\eta_{I}$ and $\eta_{II}$ increase very slowly for $s=0$ and $s=0.2$. For $s=1$, $\eta_{II}$ decreases gradually, while $\eta_{I}$ stays constant. These features are a result of the fact that an increase of $R$ can improve the accuracy of computation, but is not favorable for the enhancement of efficiency.
@@ Appendix III Approaching spin dynamics in double-excitation subspace by Gauss quadrature in real space
In Fig. A6, $P_{e}(t)$ is evaluated using the Gauss quadrature method discretization . The weight function $w(x)=\left(x/\omega_{c}\right)^{s}e^{-x/\omega_{c}}$ is chosen to construct the chain Hamiltonian and the number of lattice sites is set to $N=200$. It is evident that $P_{e}(t)$ has a similar evolution as the survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ in single-excitation subspace. However, in contrast to the CDA approach, strong fluctuations due to finite $N$ appear when $t>\sim 2$, which limits the application of Gauss quadrature in the simulation of long-term dynamics in open quantum systems.